# A local, many-worlds, model of quantum correlations with finite information flow

### v2

@@ III A model with finite information flow
Now, let us show that two “equally weighted” branches and a finite information flow are enough for simulating the outcomes of all the projective measurements $\vec{a}$ and $\vec{b}$ on a maximally entangled state. The proof is a straightful application of DH argument to Toner-Bacon model toner . The model provides a classical simulation of the outcomes by using one bit of communication between the parties and some shared randomness.
@@
Averaging on $\vec{x}_{0}$ and $\vec{x}_{1}$, it turns out that the outcomes are generated according to Eq. (2). Let us prove that. It is sufficient to show that
@@
$\begin{array}[]{l}\langle s_{A}s_{B}\rangle=\frac{1}{(4\pi)^{2}}\sum_{n_{A}=\pm 1}\int d^{2}x_{0}d^{2}x_{1}\text{sign}(\vec{a}\cdot\vec{x}_{0})\vspace{1mm}\\ \text{sign}\left[-\vec{b}\cdot(\vec{x}_{0}+n_{A}\vec{x}_{1})\right]\frac{1+n_{A}\text{sign}(\vec{a}\cdot\vec{x}_{0})(\vec{a}\cdot\vec{x}_{1})}{2}.\end{array}$ (7)
First, we note that the two terms of the sum over $n_{A}$ are identical (replace $\vec{x}_{1}$ with $-\vec{x}_{1}$ in the second term). Using the distributive property and a suitable swapping $\vec{x}_{0}\leftrightarrow\vec{x}_{1}$ , we have
@@
where $\theta(x)$ is the Heaviside function. The first term appears in the Kochen-Specker model kochen and is $-(1+\vec{a}\cdot\vec{b})/2$. Thus, Eq. (6) is proved.
Now, let us reinterpret this model within a framework à la many-worlds, where each party branches into two distinct alternatives, labeled $A_{\pm 1}$ and $B_{\pm 1}$ for Alice and Bob, respectively. The two parties are not allowed to communicate until they reach the meeting point. First, we note that the Toner-Bacon model generates the same correlation if $s_{A}$ and $s_{B}$ are both chosen with opposite sign. Let us assume that alternatives $A_{1}$ and $A_{-1}$ get the outcome $s_{A}$ in Eq. (3) and its opposite, respectively. They both generate $n_{A}$ according to Eq. (4). All these processes involve only local information that is available to Alice. A complication arises in deciding which outcomes are associated with Bob’s alternatives. Associating $s_{B}$ in Eq. (5) to alternative $B_{1}$ would require to know the value $n_{A}$, an information that is not available to Bob. If we set $n_{A}$ equal to $1$ and, finally, we pair alternative $A_{w}$ with $B_{w}$ at the meeting point, we end up making a mistake in the case $n_{A}$ were $-1$ and $\text{sign}\left[\vec{b}\cdot\vec{x}_{+}\right]\neq\text{sign}\left[\vec{b}\cdot\vec{x}_{-}\right]$, where $\vec{x}_{\pm}\equiv\vec{x}_{0}\pm\vec{x}_{1}$. Crucially, the mistake can be corrected by knowing just one classical bit of information per measurement, namely $n_{A}$ for Alice’s measurement and
@@
and $-s_{B}$, respectively. Furthermore, both Bob’s alternatives generate $n_{B}$ according to Eq. (11), which will be used to correct the pairing. Finally, Bob and Alice reach the meeting point to compare their results. Bob and Alice carry information about their outcome and the numbers $n_{A}$ and $n_{B}$, respectively. Alternatives $A_{\pm 1}$ are paired with $B_{\pm 1}$, unless $n_{A}=n_{B}=-1$, in which case the pairing is swapped. This leads to the following pairing rule
@@
From the perspective of one realization of Alice and Bob, the probability of getting outcomes $s_{A}$ and $s_{B}$ is obtained by a simple “unweighted” branch counting. First, we compute the fraction of branches getting $s_{A}$ and $s_{B}$. Then, we average over $\vec{x}_{0}$ and $\vec{x}_{1}$. Note that there is no amplitude weight of the branches.
This model is mathematically equivalent to the model in Ref. gisin , but they differ in the interpretation. The latter is a single-world model in which the parties share a nonlocal PR box as a resource for generating the outcome. In our model, we have a ”local” PR box which is used for pairing locally the alternatives at the meeting point. The statistics of the single-world model of Ref. gisin is recovered by randomly taking one of the two paired alternatives in the ”many-worlds” model. It is surprising that a simple and seemingly trivial conceptual step separates our model from that in Ref. gisin , yet it has gone unnoticed until now.
It is interesting to note that a local single-world model could be constructed in a slightly different scenario. Instead of generating the correct outcomes at the moment measurements are performed, the task shifts to generating them at the meeting point. Each party only needs to communicate two bits. However, including the whole process in this local description would lead to the paradoxical conclusion that a party’s state – and even their memory of past observations – can suddenly change. In a many-worlds framework, this macroscopic ‘rewind’ and memory reconstruction is avoided by allowing two distinct realities to evolve independently.
IV Discussion
The nonlocality arising in the framework of ontological theories can lead the proponents of Deutsch-Hayden many-worlds theory to claim the failure of a probabilistic framework. The main purpose of this paper is to show with a proof of principle that only the branching is required for a local account of the correlations of two maximally entangled qubits. There is no known proof that both determinism and the ontology of the quantum states are necessary conditions for a local theory. Furthermore, the drop of these two assumptions can lead to a more ‘economical’ theory, in which the information flow and the number of branches is made as small as possible.
Let us discuss in detail the two key differences between our model and the Deutsch-Hayden many-worlds theory – namely, the way the probabilities come out and the role played by the quantum state.
Probabilities – In the scenario simulated by the model, if Alice and Bob perform the measurements $\vec{a}$ and $\vec{b}$, then they observe outcome $s_{A}$ and $s_{B}$ with probability $(1-s_{A}s_{B}\vec{a}\cdot\vec{b})/2$. The many-worlds theory says that there are $4$ branches, one for each outcome, weighted by an amplitude whose modulus square is the observed probability. In Saunders’ terms, there is a multitude of branches such that the ratio between the number of branches with outcomes $s_{A}$ and $s_{B}$ and the number of overall branches is equal to the observed probability. Since, this probability is a real number, this multitude is actually infinite.
In our model, the story is quite different. There are only two “equally weighted” branches and, in a single run of the experiment, the number of branches in which an outcome is observed can be 00 or $1$. This implies that some outcomes are not realized in a single instance of the experiment (in MW theory, all the outcomes are realized). The observed probability of an outome is obtained by taking the fraction of branches with that outcome and averaging over different runs. This allows us to recover Born’s rule in a more natural and economical way.
@@ IV Discussion
Role of the quantum state – The second difference is more subtle. In a $\psi$-epistemic theory, the ontic state does not contain the full information about the quantum state. Rather, this information is encoded in the statistics of the ontic state. Let us consider the quantum state emerging from Alice’s measurement and how its information is encoded in Deutsch-Hayden theory and our model. In the former, the ongoing state, which has to be carried to the meeting point, is the vector $\vec{a}$ and the outcome associated with each branch. Their knowledge allows us to infer the outgoing quantum state after the measurement in each branch, meaning that the quantum state is part of the ontology. In the latter, this inference is not possible in a single run of the experiment. After the measurement, Alice generates a number $n_{A}$ and associates an outcome to each branch. This information and the knowledge of the shared random variables $\vec{x}_{0}$ and $\vec{x}_{1}$ enable us only to infer that the quantum state $\vec{a}$ in branch $A_{1}$ after the measurement is any unit vector such that Eqs. (3,4) are fulfilled. Rather, the full information about quantum state is statistically encoded in the ontic variables. This is necessary for making the information flow from the parties to the meeting point finite montina_comm_epist0 ; montina_comm_epist .
V Conclusions
@@
In conclusion, a hybrid approach that combines the randomness of ontological theories with the branching structure of many-worlds theory can address the challenges of each individual framework. On one hand, branching allows us to evade Bell’s theorem, which applies to single-world ontological theories. More broadly, it may also circumvent the contextuality requirement. For instance, a manifestation of preparation contextuality is the proof that ontological theories inherently break time symmetry leifer1 . However, with appropriate adaptations, our model provides a fully time-symmetric description of a scenario in which a qubit, in a maximally mixed state, is measured at two different times (this model will be presented in a future version of Ref. monti_erase ). The time asymmetry proved in Ref. leifer1 is transferred to the measurement devices and the subsequent comparison of results, which are inherently time-asymmetric processes. Similarly, the branching circumvents the recently proved theorem that measurements erase information monti_erase , which is another manifestation of contextuality. A many-worlds framework can also elude the Pusey-Barrett-Rudolph theorem pbr , which implies that locality is in conflict with $\psi$-epistemic theories under a single-world premise. Note that this bypass of the theorem has never been considered so far, since Deutsch-Hayden theory is local but also $\psi$-ontic. On the other hand, the introduction of randomness in the many-worlds framework can make it more natural and economical to account for the observed randomness through a simple count over a finite set of branches. Furthermore, it allows us to reduce the information flow as much as possible by stripping the quantum state of its ‘ontic’ rank. Finally, if the quantum state represents our knowledge about a system rather than an element of reality, then it is possible that not all alternatives in a superposition are necessarily realized (as occurs in our model). If there were some branch with a very faint amplitude in which Scipio was defeated in the Third Punic War, this would not necessarily imply that there is a world in which Carthage rules over Rome!
@@
Here, we have presented a proof of principle of this hybridation by showing that local measurements on two maximally entangled qubits can be locally simulated with only two ‘unweighted’ branches, finite communication and shared randomness. It is plausible that a generalization to many qubits and any quantum state will require a blow-up of the shared variables. This increase of local resources occurs also in the Deutsch-Hayden many-worlds theory deutsch , in which the local observables are represented by matrices acting on the whole Hilbert space of all involved particles. The main difference, however, is that the shared random variables in our approach would not depend on the state of the system – they would be generated before the ”game” starts and would provide a stochastic background over which the system evolves. A similar idea has been suggested in Ref. montina_non_markov within a single-world scenario. However, whether this hybrid approach can be extended to a general theory remains an open and, in our opinion, fascinating question.

### v3

@@ V A model with finite information flow
Employing the framework of Sec. III.4, let us show that two unweighted instances and a finite information flow are enough for simulating the outcomes of all the projective measurements $\vec{a}$ and $\vec{b}$ on a maximally entangled state. The proof is a simple adaptation of the Toner-Bacon model toner to the ontological model of Sec. III.4 with $m=2$ instances per party. The Toner-Bacon model provides a classical simulation of the outcomes by using one bit of communication between the parties and some shared randomness. Let us first review this single-world model.
@@
Averaging on $\vec{x}_{0}$ and $\vec{x}_{1}$, it turns out that the outcomes are generated according to Eq. (11). Let us reproduce the proof of Ref. toner by using the Kochen-Specker model kochen as shortcut in the last step. It is sufficient to show that
@@
$\begin{array}[]{l}\langle s_{A}s_{B}\rangle=\frac{1}{(4\pi)^{2}}\sum_{n_{A}=\pm 1}\int d^{2}x_{0}d^{2}x_{1}\text{sign}(\vec{a}\cdot\vec{x}_{0})\vskip 2.84526pt\\ \text{sign}\left[-\vec{b}\cdot(\vec{x}_{0}+n_{A}\vec{x}_{1})\right]\frac{1+n_{A}\text{sign}(\vec{a}\cdot\vec{x}_{0})(\vec{a}\cdot\vec{x}_{1})}{2}.\end{array}$ (18)
First, we note that the two terms of the sum over $n_{A}$ are identical (replace $\vec{x}_{1}$ with $\vec{x}_{1}$ in the second term). Using the distributive property and a suitable swapping $\vec{x}_{0}\leftrightarrow\vec{x}_{1}$ , we have
@@
where $\theta(x)$ is the Heaviside function. The first term appears in the Kochen-Specker model kochen and is $(1+\vec{a}\cdot\vec{b})/2$. Thus, Eq. (17) is proved.
Now, let us reinterpret this model within the framework of Sec. III.4 with $m=2$ instances per party. The resulting model is schematically represented in Fig. 1. The variables $\eta_{A}$ and $\eta_{B}$ in Sec. III.4 contain the integers $n_{A}$ and $n_{B}$, respectively, and a copy of the random variables $\vec{x}_{0}$ and $\vec{x}_{1}$. Initially, the two instances are indistinguishable. When the measurements are performed, each party branches into two distinct alternatives, labeled $A_{\pm 1}$ and $B_{\pm 1}$ for Alice and Bob, respectively. The two parties are not allowed to communicate until they reach the meeting point, where each alternative of Alice is paired to one alternative of Bob according to some rule that we will determine. First, we note that the Toner-Bacon model generates the same correlation if $s_{A}$ and $s_{B}$ are both chosen with opposite sign. Let us assume that alternatives $A_{1}$ and $A_{-1}$ get the outcome $s_{A}$ in Eq. (14) and its opposite, respectively. They both set $n_{A}$ according to Eq. (15). All these processes involve only local information that is available to Alice. Thus, the ontic state $\lambda_{A}$ of Eq. (7) evolves locally according to Eq. (8). A complication arises in deciding which outcomes are associated with Bob’s alternatives. Associating $s_{B}$ in Eq. (16) to alternative $B_{1}$ would require to know the value $n_{A}$, an information that is not available to Bob. If we set $n_{A}$ equal to $1$ and, finally, we pair alternative $A_{w}$ with $B_{w}$ at the meeting point, we end up making a mistake in the case $n_{A}$ were 1-1 and $\text{sign}\left[\vec{b}\cdot\vec{x}_{+}\right]\neq\text{sign}\left[\vec{b}\cdot\vec{x}_{-}\right]$, where $\vec{x}_{\pm}\equiv\vec{x}_{0}\pm\vec{x}_{1}$. Crucially, the mistake can be corrected during the pairing by knowing just one classical bit of information per measurement, namely $n_{A}$ for Alice’s measurement and
@@
and $s_{B}$, respectively. Furthermore, both Bob’s alternatives generate $n_{B}$ according to Eq. (22), which will be used to define the pairing. Also the ontic state $\lambda_{B}$ of Eq. (7) evolves locally according to Eq. (8). Thus, the overall model is local. Finally, Bob and Alice reach the meeting point to compare their results. Bob and Alice carry information about their outcome and the numbers $n_{A}$ and $n_{B}$, respectively. Alternatives $A_{\pm 1}$ are paired with $B_{\pm 1}$, unless $n_{A}=n_{B}=-1$, in which case the pairing is swapped. This leads to the following pairing rule
@@
Let us stress that there are only two instances after the pairing. If $(n_{A},n_{B})\neq(-1,-1)$, then Alice and Bob agree that the pair of outcomes is $(s_{A},s_{B})$ in one instance and $(-s_{A},-s_{B})$ in the other instance. Conversely, if $(n_{A},n_{B})=(-1,-1)$, the pairing is flipped and the instances are $(s_{A},-s_{B})$ and $(-s_{A},s_{B})$. It is worth to note that only $2$ of the $4$ possible combinations of outcomes are realized in each run. If there were $4$ unweighted instances deterministically realizing all the combinations, the probability of observing one pair of outcomes would be $1/4$ in each run. To match quantum theory using a finite number of unweighted instances, the final pair of outcomes must be stochastic – that is, they must vary from run to run. From the perspective of one realization of Alice and Bob, the probability of getting a pair of outcomes $(s^{\prime},s^{\prime\prime})\in\{-1,1\}^{2}$ is obtained by a simple count of unweighted instances. First, we compute the fraction of instances getting $(s^{\prime},s^{\prime\prime})$. Then, we average over $\vec{x}_{0}$ and $\vec{x}_{1}$. There is no amplitude weight of the branches.
@@
Figure 1: Local model simulating correlations of local projective measurements on two maximally entangled qubits. Alice and Bob receive each two random vectors $\vec{x}_{0}$ and $\vec{x}_{1}$ from a common source. Given projective measurements $\vec{a}$ and $\vec{b}$, they locally generate $\{n_{A},\vec{s}_{A}\equiv(s_{A},-s_{A})\}$ and $\{n_{B},\vec{s}_{B}\equiv(s_{B},-s_{B})\}$, respectively, by using only locally available information. The numbers $s_{A}$, $n_{A}$, $s_{B}$ and $n_{B}$ are generated according to Eqs. (14152322), respectively. Finally, they meet at a common point and each instance of one party is paired with one instance of the other party according to the rule in the figure. This results in two instances of the overall outcomes.
Summarizing, the model is as follows.
The $\psi$-epistemic local model.
Simulation of local von Neumann measurements with two instances and one communicated bit per party in Bell’s scenario of two maximally entangled qubits (Fig. 1).
  • Alice and Bob receive the random vectors $\vec{x}_{0}$ and $\vec{x}_{1}$.
  • In instance $A_{\pm 1}$, Alice generates outcome $\pm s_{A}$ according to Eq. (14), given the measurement setting $\vec{a}$. She also sets $n_{A}$ by using Eq. (15).
  • In instance $B_{\pm 1}$, Bob generates outcome $\pm s_{B}$ according to Eq. (23), given the measurement setting $\vec{b}$. He sets $n_{B}$ by using Eq. (22).
  • Finally, they reach the meeting point. Instance $A_{\pm 1}$ is paired with instance $B_{\pm 1}$ if $(n_{A},n_{B})\neq(-1,-1)$, otherwise the pairing is swapped.
At each run of the simulation, two instances of the overall outcomes are generated. The quantum probabilities are obtained by selecting one of the two instances at random with equal probability (assumption of equiprobability of instances in Sec. III.1).
It is interesting to note that the pairing depends only on $n_{A}$ and $n_{B}$. There is no need to carry the initial random vectors $\vec{x}_{0}$ and $\vec{x}_{1}$, which are used only for generating $s_{A}$, $s_{B}$, $n_{A}$ and $n_{B}$.
This model is mathematically equivalent to the model in Ref. gisin , but they differ in the interpretation. The latter is a single-world model in which the parties share a nonlocal PR box as a resource for generating the outcome. In our model, we have a “local” PR box which is used for pairing locally the alternatives at the meeting point. The statistics of the single-world model of Ref. gisin is recovered by randomly taking one of the two paired alternatives in the two-instances model. It is surprising that a simple and seemingly trivial conceptual step separates our model from that in Ref. gisin , yet it has gone unnoticed until now.
It is interesting to note that a local single-world model could be constructed in a slightly different scenario. Instead of generating the correct outcomes at the moment measurements are performed, the task shifts to generating them at the meeting point. Each party only needs to communicate two bits. However, including the whole process in this local description would lead to the paradoxical conclusion that a party’s state – and even their memory of past observations – can suddenly change. In a many-instances framework, this macroscopic “rewind” and memory reconstruction is avoided by allowing two distinct realities to evolve independently.
VI Discussion
The nonlocality arising in the framework of single-world ontological theories can lead the proponents of Deutsch-Hayden many-worlds theory to claim the failure of a probabilistic framework in which $\psi$ represents a mere state of knowledge. The main purpose of this paper is to show with a proof of principle that the coexistence of multiple realities is sufficient for a local account of the correlations of two maximally entangled qubits. There is no known proof that both determinism and the ontology of the quantum states are necessary conditions for a local theory. Furthermore, the drop of these two assumptions can lead to a more “economical” theory, in which the information flow is made as small as possible and the quantum probabilities are obtained by a simple count of unweighted instances.
Let us discuss in detail the two key differences between the model in Sec. V and the Deutsch-Hayden many-worlds theory – namely, the way the probabilities come out and the role played by the quantum state.
Probabilities – In the scenario simulated by the model, if Alice and Bob perform the measurements $\vec{a}$ and $\vec{b}$, then they observe outcome $s_{A}$ and $s_{B}$ with probability $(1-s_{A}s_{B}\vec{a}\cdot\vec{b})/4$. The many-worlds theory says that there are $4$ branches, one for each outcome, weighted by an amplitude whose modulus square is the observed probability. In Saunders’ terms, there is a multitude of branches of microstates such that the ratio between the number of branches with outcomes $s_{A}$ and $s_{B}$ and the number of overall branches is equal to the observed probability. Since this probability is a real number, this multitude is actually infinite.
In our model, the story is quite different. There is a finite number of unweighted instances (namely, two instances). The observed probability of the final pair of outcomes $(s_{A},s_{B})$ is obtained by taking the fraction of instances with outcome $(s_{A},s_{B})$ (the fraction is 0 or $1/2$ in the specific model) and averaging over different runs. Like in the general framework of Sec. III, the different instances have equal probability of being experienced by an observer. Note that some outcomes are not realized in a single run of the experiment (in MW theory, all the outcomes are realized in each run). Indeed, two instances cannot realize all the $4$ possible combinations of outcomes. The remarkable point of the model is not the reduction from $4$ to $2$ realized branches, but the equiprobabilities of the two instances. This is possible because of the stochasticity of the model. As previously said, the derivation of the branch weights of many-worlds theory from a branch counting à la Saunders saunders would require a division of each of the $4$ branches into an infinite multitude of equally probable instances (see also Sec. III.2.1, in which it is shown that determinism is recovered in the limit of infinite instances in our framework).
Recalling the general discussion of Sec. III, let us transliterate this feature of the model to an experiment generating the superposition
$\psi_{+}|+\rangle+\psi_{-}|$-$\rangle$ (25)
where $|\pm\rangle$ are two macroscopically distinct states, and $|\psi_{-}|^{2}=10^{-40}|\psi_{+}|^{2}$. In many-worlds theory, both the alternatives are realized, however we expect to experience $|+\rangle$ with near-certainty. In our framework, we have a finite number $m$ of instances such that the ratio between the number of instances associated with state $|$-$\rangle$ and the number of all the instances is equal to our expectation of experiencing $|$-$\rangle$ in a single run of the experiment. If the total number of instances is small ($m\ll 10^{40}$), this ratio is zero in most of the runs. That is, the branch $|$-$\rangle$ is almost never realized. The ratio turns out to be equal to $|\psi_{-}|^{2}$ only on average over a large number of runs (like in any single-world ontological theory). This point of view is in the middle between a single-world ontological theory, such as de Broglie-Bohm theory, and many-worlds theory. In the former, there is some ontic variable saying that the system is always realized in one of the two superposed states. The latter says that both the states are always realized in parallel worlds. As a single-world theory is nonlocal, as proved by Bell, we may need more than one world for a local description, but between 11 (single-world ontological theories) and infinity (many-worlds theory plus Saunders’ branch counting) there is any integer greater than 11!
@@ VI Discussion
Role of the quantum state – The second difference is more subtle. In a $\psi$-epistemic theory, the ontic state does not contain the full information about the quantum state. Rather, this information is encoded in the statistics of the ontic state. Let us consider the quantum state emerging from Alice’s measurement and how its information is encoded in Deutsch-Hayden theory and our model. In the former, the ongoing state, which has to be carried to the meeting point, is the vector $\vec{a}$ and the outcome associated with each branch. Their knowledge allows us to infer the outgoing quantum state after the measurement in each branch, meaning that the quantum state is part of the ontology. In our model, this inference is not possible in a single run of the experiment. After the measurement, Alice generates a number $n_{A}$ and associates an outcome to each branch. This information and the knowledge of the shared random variables $\vec{x}_{0}$ and $\vec{x}_{1}$ enable us only to infer that the quantum state $\vec{a}$ in branch $A_{1}$ after the measurement is any unit vector such that Eqs. (14,15) are fulfilled. Rather, the full information about quantum state is statistically encoded in the ontic variables. This is necessary for making the information flow from the parties to the meeting point finite montina_comm_epist0 ; montina_comm_epist . It is worth to recall again that the whole ontic state after the measurements is not actually required at the meeting point, since $\vec{x}_{0}$ and $\vec{x}_{1}$ are not involved in the pairing rule.
VII Conclusions
In conclusion, we have introduced a general probabilistic framework that encompasses both single-world and many-worlds theories as limiting cases. This hybrid approach, which combines the randomness of single-world ontological theories with the multiplicity of coexisting realities found in many-worlds theory, offers a way to address the challenges inherent in each individual previous framework.
@@ VII Conclusions
  • On one hand, the coexistence of diverse realities allows us to evade Bell’s theorem, which applies to single-world ontological theories. More broadly, it may also circumvent the contextuality requirement. For instance, a manifestation of preparation contextuality is the proof that ontological theories inherently break time symmetry leifer1 . However, with appropriate adaptations, our model provides a fully time-symmetric description of a scenario in which a qubit, in a maximally mixed state, is measured at two different times monti_erase . The time asymmetry proved in Ref. leifer1 is transferred to the measurement devices and the subsequent comparison of results, which are inherently time-asymmetric processes. Similarly, the branching circumvents the recently proved theorem that measurements erase information monti_erase , which is another manifestation of contextuality. An open question is whether a many-instances framework can also elude the Pusey-Barrett-Rudolph (PBR) theorem pbr , which implies that quantum mechanics is in conflict with $\psi$-epistemic theories under a single-world premise and an assumption of preparation independence (a kind of locality condition). Note that this bypass of the theorem has never been considered so far, since Deutsch-Hayden theory is local but also $\psi$-ontic. It is also interesting to note that the hypothesis of preparation independence may be questioned in the framework of nonlocal single-world ontological theories, but it becomes mandatory in our framework, which aims to provide a locally causal structure. One possible way to circumvent the theorem is to assume that multiple preparations may coexist, which is more drastic than assuming that multiple outcomes coexist.
  • On the other hand, the randomness of the framework makes it possible to recover Born’s rule from a simple count over a finite set of instances. Furthermore, it can lead to a reduction of the information flow as much as possible by stripping the quantum state of its “ontic” rank. Finally, if the quantum state represents our knowledge about a system rather than an element of reality, then it is possible that not all alternatives in a superposition are necessarily realized (as discussed in Sec. III.1 and illustrated by our model). If, according to prior knowledge, there existed a branch with a very faint amplitude in which Scipio was defeated in the Third Punic War, this would not necessarily imply that there is a world where Carthage rules over Rome. That is, the multiple instances of the framework discussed in Sec. III.1 may not realize all possible branches.
@@
Here, we have presented a proof of principle of this hybridization by showing that local measurements on two maximally entangled qubits can be locally simulated with only two unweighted instances, finite communication and shared randomness. It is plausible that a generalization to many qubits and any quantum state will require a blow-up of the shared variables. This increase of local resources occurs also in the Deutsch-Hayden many-worlds theory deutsch , in which the descriptors are represented by matrices acting on the whole Hilbert space of all involved particles. The main difference, however, is that the shared random variables in our approach would not depend on the state of the system – they would be generated before the “game” starts and would provide a stochastic background over which the system evolves. A similar idea has been suggested in Ref. montina_non_markov within a single-world scenario. Our purpose has been to introduce a general framework with a simple proof of principle, rather than a comprehensive theory. Whether this hybrid approach can be extended to a general local theory distinct from DH theory remains an open and, in our opinion, fascinating question. A negative answer to this question would suggest that every locally causal theory of quantum mechanics is essentially the DH theory.
A generalization of the presented model to higher dimensional systems can be obtained from the approximate model in Ref. montina_approx which extends the Toner-Bacon protocol to arbitrary dimensions. Although this generalized model is not exact, its error is very small (about $1\%$) for few qubits per party. Its adaptation to our many-instances framework would provide a very accurate model for a sufficiently low dimension. This adaptation can suggest improvements for reducing the error or finding exact protocols.
From a broader perspective, this framework provides a generalized setting for quantum communication complexity. Rather than just quantifying the communication cost of classically simulating quantum entanglement, one can consider a scenario involving a third party – a referee – who receives multiple alternative outcomes and determines how to pair them. In this setting, the question concerns not only the minimal communication cost from the two parties to the referee, but also the minimal number of parallel unweighted instances required for a local simulation of quantum correlations. The standard communication complexity problem of quantifying the required communication cost for simulating entanglement is a subclass of this generalized setting with the slight variation that the generation of one of the outcomes is postponed and delegated to the referee. The subclass corresponds to the situation in which there is one instance and one of the two parties – say Bob – communicates to the referee the full information about his choice. Let us underline that, in this case, Bob is not supposed to generate the outcome, but delegates the task to the referee. A further generalization could allow for multiple parallel choices of measurements performed by Alice and Bob. Focusing only on one side of the problem, one can consider $\psi$-ontic models with infinite communication and find the minimal number of instances for an exact local simulation. It can be interesting, for example, to investigate some kind of ‘pilot’ theory with a finite number of instances in which each instance is locally piloted by Deutsch-Hayden descriptors (rather than the wave-function like in pilot-wave theories). The task would be to find the minimal number of instances and the pairing rules that lead to an exact simulation of quantum correlations for more general entangled states.
In the context of this generalized setting in quantum communication complexity, there is an interesting aspect that is worth mentioning and offers perspectives for future research. On one hand, it is known that the communication cost of simulating quantum correlations goes to zero as the entangled state approaches a product state renner . Thus, if we adopt the communication cost as a measure of ‘nonlocality’, weakly entangled states are less ‘nonlocal’ than strongly entangled states (here, ‘nonlocality’ refers to single-world scenarios). On the other hand, maximally entangled states can be simulated using a single ‘nonlocal’ PR-box gisin , whereas weakly entangled states require more than one PR-box brunner . Thus, adopting the number of PR-boxes as a measure of nonlocality, we obtain the opposite conclusion that weakly entangled states are more ‘nonlocal’ than maximally entangled states. In our generalized setting of communication complexity, the task is not just to evaluate the minimal communication cost, but also the minimal number of parallel realities (instances) for a local account of quantum correlations. It is entirely possible that these two quantities behave differently as functions of the ‘degree’ of entanglement. In particular, the number of required instances is likely related to the second measure of ‘nonlocality’ (with PR-boxes). Thus, a local model with multiple instances could potentially ‘explain’ why weakly entangled states are more robust against the detection loophole in an experimental Bell test loophole . Indeed, if more instances are required for a correct simulation, a single-world protocol is less likely to succeed by discarding some simulation runs instead of correctly pairing the instances. Thus, our framework is not just relevant in quantum foundations, but suggests new technical problems which have their own appeal for classical simulations of quantum entanglement and quantum communication. Our framework may provide new insights into some of the counter-intuitive behaviors of entanglement.
Another line of future research can be inspired by the approach of many interacting worlds in Ref. hall , which has been built on ground of the dBB theory. As in our framework, this approach employs a finite number of worlds but is equivalent to quantum mechanics only in the limit of infinite worlds. Thus, it is experimentally testable. The approach is nonlocal, so that one can wonder whether a modification making it “local” is possible. For example, in the search of such a modified approach, one could start from the descriptors, rather than the wave function.